# Shaping Sparse Rewards in Reinforcement Learning: A Semi-supervised Approach

### v3

@@ Abstract
In many real-world scenarios, reward signal for agents are exceedingly sparse, making it challenging to learn an effective reward function for reward shaping. To address this issue, the proposed approach in this paper performs reward shaping not only by utilizing non-zero-reward transitions but also by employing the Semi-Supervised Learning (SSL) technique combined with a novel data augmentation to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving the efficacy of reward shaping. Experimental results in Atari and robotic manipulation demonstrate that our method outperforms supervised-based approaches in reward inference, leading to higher agent scores. Notably, in more sparse-reward environments, our method achieves up to twice the peak scores compared to supervised baselines. The proposed double entropy data augmentation enhances performance, showcasing a 15.8% increase in best score over other augmentation methods
Introduction
@@
The sparse reward problem is a core challenge in Reinforcement Learning (RL) when solving real-world tasks (Kober, Bagnell, and Peters 2013). In supervised learning, supervision signals are provided by the training data. In reinforcement learning, rewards take on the role of supervision signals, guiding the agent to optimize its policy. Many real-world tasks naturally have the feature of delayed or infrequent rewards due to the complexity and nature of the tasks. In tasks like Go (Silver et al. 2016), navigation (Wang et al. 2020), or robotic arm manipulation (Gu et al. 2017; Riedmiller et al. 2018), agents only receive rewards upon successfully achieving the final goal, such as winning a game, reaching a target, or completing a grasp, with no feedback for intermediate steps. Sparse reward provides little immediate feedback to guide the agent’s exploration and high variance in returns, makes it difficult to learn the optimal policy (Plappert et al. 2018).
A potential solution to tackle the sparse reward problem is reward design and learning (Ng, Harada, and Russell 1999). Methods in this category often operate on the replay buffer, processing off-policy data to implement their respective reward design (Schaul et al. 2015; Andrychowicz et al. 2017; Peng et al. 2019). Pathak et al. (2017) propose the Intrinsic Curiosity Module (ICM) that formulates curiosity as the error in an agent’s ability to predict the consequence of its own actions learned by a self-supervised inverse dynamics model. However, curiosity-driven methods face a significant challenge: curiosity, as an intrinsic reward, is inherently decoupled from the actual objectives and tasks. This may cause agents to overly focus on “new” but meaningless states. Kumar, Peng, and Levine (2019) propose Reward Conditioned Policy (RCP) that uses supervised learning as an incentive, thereby eliminating reliance on manually designed novelty as the objective. Building upon this, Self-Supervised Online Reward Shaping (SORS, Memarian et al. (2021)) formalizes a self-supervised learning framework and apply it to sparse-reward environments. In SORS, the original sparse reward provides a self-supervisory signal for reward inference by ranking trajectories that the agent observes, while the policy update is performed with the newly inferred, typically dense reward function. However, methods based on non-expert trajectories under supervised learning suffer from poor sample efficiency (Peng et al. 2019), especially when non-zero-reward transitions are extremely rare in sparse reward cases.
@@
In order to bridge the gap in handling sparse rewards, our approach performs reward shaping not only by utilizing non-zero-reward transitions but also by employing Semi-Supervised Learning (SSL) technique to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving reward shaping. Explicitly, we propose an RL algorithm that applies the idea of SSL, called Semi-Supervised Reward Shaping framework (SSRS), and also propose a new data augmentation technique called double entropy data augmentation applied on non-image transitions as the perturbation method for SSL. By progressively approximating the state-value function and the action-state function in the reward estimator, we shape the reward in sparse-reward trajectories. Consistency regularization to both zero-reward and non-zero-reward transitions is adopted for the optimization of the reward estimator. Additionally, a monotonicity constraint is incorporated over the two components of the reward shaping estimator to further reduce the discrepancy between the shaped reward distribution and the true reward distribution.
We evaluate the performance of SSRS in reward-sparse Atari and robotic manipulation environments, comparing it with SORS (Memarian et al. 2021), RCP (Kumar, Peng, and Levine 2019) methods. Our model demonstrates performance superior than other supervised-based approaches and achieves a maximum increase over RCP by a factor of two in reaching higher best scores. We further validate that the proposed double entropy data augmentation enhances performance showcasing a 15.8% increase in best score compared to other augmentation methods.
Related Work
@@
There are different reward shaping methods for different objectives: Count-based methods (Choi et al. 2019; Ostrovski et al. 2017; Bellemare et al. 2016) incentivize agents based on the rarity of states, while curiosity-driven exploration methods reward the agent’s exploratory behavior. ICM Pathak et al. (2017) also pursues rarity, but operationalizes it by defining curiosity as the error in the agent’s ability to predict the consequences of its own actions in a visual feature space. This curiosity signal serves as an intrinsic reward, driving the agent to explore novel states. Andrychowicz et al. (2017) propose Hindsight Experience Replay (HER) that enables learning from failures by treating unachieved goals as alternative goals for experience replay.
Approaches using supervised learning for reward shaping don’t rely on specific objectives. The study (Kumar, Peng, and Levine 2019) employs supervised learning techniques, viewing non-expert trajectories collected from sub-optimal policies as optimal supervision to do the reward shaping. Memarian et al. (2021) leverage self-supervised methods to extract signals from trajectories while simultaneously updating the policy.
@@
We extend supervised learning to a semi-supervised paradigm and introduce SSL techniques into the RL framework, specifically employing consistency regularization. and pseudo-labeling (Bachman, Alsharif, and Precup 2014) in the trajectory space. In the realm of SSL, consistency regularization leverages unlabeled data by capitalizing on the invariance assumption—that the model should yield consistent predictions when presented with perturbed variants of identical inputs. Berthelot et al. (2019); Sohn et al. (2020) further improve this methodology on image classification tasks. Pseudo-labeling (Lee 2013) utilizes the model’s predicted hard labels (obtained via $\arg\max$ operations) for unlabeled data, retaining only those predictions whose confidence scores $\mathbf{q}$ exceed the threshold $\tau$, and optimizes them via:
$\mathcal{L}_{u}=\mathds{1}(\max(\mathbf{q})\geq\tau)\cdot\mathcal{H}(\mathbf{q},\arg\max(\mathbf{q})).$ (1)
Note we directly write $\mathbf{q}$, omitting the dependence of $\mathbf{q}$ on the parameters and $\mathds{1}(\cdot)$ is the indicator function.
Another related direction is the application of data augmentation for trajectories (Shorten and Khoshgoftaar 2019) in RL algorithms (Yarats, Kostrikov, and Fergus 2021). Most of them process image data to enhance the agent’s generalization performance. Hansen and Wang (2020) decouple augmentation from policy learning and is found to significantly advance sample efficiency, generalization, and stability in training over vision-based RL methods. Yarats et al. (2022) use data augmentation to learn directly from pixels and is able to solve complex humanoid locomotion tasks directly from pixel observations. Studies like Lin et al. (2019) generate feasible trajectories based on symmetries observed in the trajectory space for robot control tasks, thereby constructing an Invariant Transform Experience Replay framework to address the issue of high sample requirements.
Methodology
@@
Figure 1: The figure illustrates how the SSRS method leverages both non-zero reward transitions and sparse reward transitions for reward shaping through a semi-supervised learning (SSL) approach incorporating strong-weak data augmentation. The augmentation techniques include Cutout, Smooth, and Double Entropy Augmentation, among others (see Table 2 in the Appendix for details).
@@
In this section, we introduce the proposed method and its pipeline is graphically illustrated in Figure 1, which consists of two components: a reward estimator optimized through semi-supervised learning, and a reward estimator optimized through semi-supervised learning and the data augmentation technique for zero-reward transitions that serves as the perturbation required for semi-supervised learning. We will elaborate the key parts of the method in the following subsections and the detailed steps are shown in Algoirthm 1.
Semi-Supervised Reward Estimator
@@
The reward estimator is the very module that takes (augmented) transitions as input and outputs estimated rewards (see Figure 1). During the training phase, the estimator is optimized via consistency regularization using the cross-entropy. In the prediction phase, the final estimated reward is generated through $\arg\max$ selection, choosing the $z$ value corresponding to the peak of the confidence score distribution. The SSRS framework proposed in this paper, aims to fit an optimal state-value function $V^{*}(s)$ and an optimal action-value function $Q^{*}(s,a)$, which are used to shape the rewards of trajectory. Note we use $Q(s,a)$ and $V(s)$ to denote aforementioned function before parameter optimization. Given an one-step trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, the estimated reward can be calculated from the estimations of the state-value function $V(s)$ and the action-value function $Q(s,a)$ as follows. We first define the confidence score vector of timestep $t$ over a fixed reward set $Z=\{z_{i},\ i=1,\dots,N_{z}\}$ in Eq.(2), where $N_{z}$ is a hyperparameter controlling the number of estimated rewards. The value of fixed reward set is initialized according to the collected true reward value, and is updated throughout the process of agent’s interactions with environment.
By the following Eq.(2), we can get confidence vector $\mathbf{q_{t}}\in\mathbb{R}^{N_{z}}$ at timestep $t$, with each element $q_{i}\in\mathbf{q_{t}}$ corresponding to a estimated reward $z_{i}\in Z$, computed as
$\mathbf{q_{t}}=\beta Q(s_{t},a_{t})+(1-\beta)V(s_{t+1}).$ (2)
The reward $z$ estimated from the trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$ essentially evaluates the quality of the agent’s interaction with the environment at the current step based on the trajectory. We utilize $Q(s_{t},a_{t})$ to assess the long-term return (i.e., the cumulative future reward) of taking action $a_{t}$ in state $s_{t}$, and combine this with the future cumulative return estimation $V(s_{t+1})$ of transitioning to state $s_{t+1}$ after taking action $a_{t}$ in state $s_{t}$, which gives us the estimated $z$ in the form of linear combination of the two, with $\beta>0$. The estimate reward value $z_{t}$ is selected with maximum confidence score above the threshold $\lambda$, i.e., $z_{t}=\alpha(\mathbf{q_{t}},\lambda)$, where
@@
Note that we will use $\mathbf{q}_{t}(\theta)$ in subsequent sections to explicitly denote its dependence on $\theta$, which consists of $\theta_{1}$ (the parameter of the $Q$ function) and $\theta_{2}$ (the parameter of the $V$ function). During the prediction stage, we apply shaping to a proportion $p_{u}$ of transitions in the buffer, where $p_{u}$ dynamically adjusts throughout training. Specifically, $p_{u}$ is initially set proportional to $\log(\mathcal{N}_{r})$, then increased to match $\mathcal{N}_{r}$ during the mid-training phases, before being reduced again to scale with $\log(\mathcal{N}_{r})$. Here, $\mathcal{N}_{r}$ represents the count of trajectories containing genuine reward signals. This prevents excessive variance in the initial policy gradient updates or $Q$-network updates, while in later stages, phases out reliance on shaped rewards as the policy converges toward optimal behavior.
@@
In order to obtain the optimal $\theta$ value for accurately estimating the reward value of each trajectory $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, SSRS framework first optimizes the following loss function using only transitions with non-zero rewards, so as to minimize the discrepancy between the reward estimator’s predictions and the ground truth values for these transitions. The handling of zero-reward transitions will be addressed in consistency regularization section.
@@
$L_{r}=\frac{1}{\mu B}\sum_{t=1}^{\mu B}\mathds{1}(\max(\mathbf{q}_{t}(\theta))\geq\lambda)\left({r_{t}}-\alpha(\mathbf{q_{t}}(\theta),\lambda)\right)^{2}.$ (3)
Note that $\mu$ represents the proportion of non-zero rewards in the buffer, which quantifies the sparsity level, and $B$ is the batch size of trajectories.
Monotonicity Constraint
Since the estimator employs two neural networks, we introduce a constraint to regulate their outputs and stabilize the training process. The value function $V(s_{t})$ provides a global baseline for a state, while advantage function captures the relative advantage or disadvantage of action $a$ with respect to this baseline. This separation makes it easier to assess the relative importance of specific actions, and thus compensates for the insensitivity to actions because the disturbance of data augmentation only takse place on state $s$. By utilizing the relationship between $V$ and $Q$, we can introduce the monotonicity constraint in the SSRS framework as follows, which quantifies the mean square positive advantage function values. Define advantage function $\delta_{t}(\theta)=Q(s_{t},a_{t},\theta_{1})-V(s_{t},\theta_{2})$ where $\theta$ is the combination of $\theta_{1}$ and $\theta_{2}$, we have
$L_{QV}=\begin{cases}0,&\delta_{t}(\theta)<0,\\ \displaystyle\frac{1}{\mu B}\sum_{t=1}^{\mu B}\left(\delta_{t}(\theta)\right)^{2},&\text{else}.\end{cases}$ (4)
Minimizing the objective loss $L_{QV}$ over the parameter $\theta$ of can achieve a more stable update, which will be validated in experiment section, and is crucial to SSRS framework’s success.
Consistency Regularization
Consistency regularization is a paradigm in the field of SSL (Sajjadi, Javanmardi, and Tasdizen 2016), and is also the key to optimize the reward estimator with the limited feedback reward signal in the sparse reward trajectories. The reward estimator, as a crucial component in SSRS, needs to capture the invariance of trajectories, which refers to the estimated reward unaffected by minor variations in input transitions, aligning with the characteristics of consistency regularization. Furthermore, in scenarios with sparse rewards, applying weak strong augmentations to transitions with zero and non-zero reward transitions before conducting consistency regularization further enhances the generalization ability of the reward estimator in sparse reward environments.
The weak and strong augmentations refer to Table 2 in Appendix, where we consider Gaussian noise with smaller parameters as weak augmentation, and other augmentation such as smoothing, translation, and cutout, with larger parameters as strong augmentation. In conventional data augmentation for weak-strong augmentation, standard techniques such as cutout, flipping, and rotation are typically employed. The unitless property of Entropy makes it particularly useful in data augmentation, as it allows for a fair comparison of information content across different features, regardless of their scale or domain. Entropy is essentially a measure of uncertainty or randomness in a probability distribution, and it is calculated in a way that normalizes the result, making it dimensionless. Building on this intuition, we further adapt a novel double entropy data augmentation method to serve in SSRS on non-image tasks. The implementation details are provided in Appendix.
@@
More specifically, for transitions with zero rewards, the reward estimator computes the loss as shown in Eq.(5), which is the loss between the strong augmentation term and weak augmentation term following typical loss design in consistency regularization. Assuming continuity of trajectories in the metric space, it calculates the confidence of reward values after weak and strong augmentations, denoted by $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$, respectively. The calculation of $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$ follows Eq.(2), simply replacing the input state s with its weakly or strongly augmented version. Here we omit their dependence on $\theta$ for simplicity. For confidence values greater than the threshold in weak augmentation, the one-hot operation is performed (Bachman, Alsharif, and Precup 2014), denoted as $\mathbf{q_{t}^{w,\,\prime}}$, and then used to compute the cross-entropy loss with the normalized confidence values greater than the threshold in strong augmentation, denoted as $\mathbf{q_{t}^{s,\,\prime}}$. The loss function is denoted as.
$L_{s}=\frac{1}{(1-\mu)B}\sum_{t=1}^{(1-\mu)B}$ $[\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$
$\cdot\mathcal{H}\left(\mathbf{q_{t}^{w,\,\prime}},\,\mathbf{q_{t}^{s,\,\prime}}\right)],$ (5)
@@
The algorithm for value-based SSRS with synchronous update—i.e., value function update and reward shaping carry out simultaneously— is outlined in Algorithm 1. The agent begins by collecting experience transitions. The reward estimation network is then initialized through back-propagation, which can be performed either synchronously or asynchronously using the loss function. The complete SSRS loss function, as specified in Eqs. (3), (4) and (5), combines these components through the weighting parameter $\alpha$. In Appendix, we provide the derivations of the “approximate” gradients of loss functions $L_{QV}$, $L_{r}$ and $L_{s}$ with respect to $\theta$ in Appendix, after approximating the indicator functions by sigmoid function.

### v4

@@ Abstract
In many real-world scenarios, reward signal for agents are exceedingly sparse, making it challenging to learn an effective reward function for reward shaping. To address this issue, the proposed approach in this paper performs reward shaping not only by utilizing non-zero-reward transitions but also by employing the Semi-Supervised Learning (SSL) technique combined with a novel data augmentation to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving the efficacy of reward shaping. Experimental results in Atari and robotic manipulation demonstrate that our method outperforms supervised-based approaches in reward inference, leading to higher agent scores. Notably, in more sparse-reward environments, our method achieves up to twice the peak scores compared to supervised baselines. The proposed entropy augmentation enhances performance, showcasing a 15.8% increase in best score over other augmentation methods.
1 Introduction
@@
The sparse reward problem is a core challenge in Reinforcement Learning (RL) when solving real-world tasks Kober et al. (2013). Many real-world tasks naturally have the feature of delayed or infrequent rewards due to the complexity and nature of the tasks. Such sparse reward signals provide minimal immediate feedback to guide the agent’s exploration, while also inducing high variance in returns. These properties make it challenging to learn an optimal policy effectively Plappert et al. (2018).
In general, densifying reward with reward design or reward shaping Ng et al. (1999) methods face the inescapable predicament of reward hacking Hadfield-Menell et al. (2017). This phenomenon can lead to a shift in the optimal policy, or cause the agent to learn a behavior that deviates from the intended design. In response to this dilemma, supervised-based reward shaping methods demonstrate greater fidelity to the original reward formulation. Moreover, Memarian et al. (2021) prove that, under mild assumptions, specific modifications to the reward function can still preserve the same set of optimal policies. However, methods based on non-expert demonstrations under supervised learning suffer from poor sample efficiency Peng et al. (2019), especially when non-zero-reward transitions are extremely rare in sparse reward cases.
@@
In order to bridge the gap in handling sparse rewards, our approach performs reward shaping not only by utilizing non-zero-reward transitions but also by employing Semi-Supervised Learning (SSL) techniques to learn trajectory space representations from the majority of transitions, i.e., zero-reward transitions, thereby improving reward shaping. Explicitly, we propose an RL algorithm that applies the idea of SSL, called Semi-Supervised Reward Shaping framework (SSRS), which alternates between policy updates using a base RL algorithm and transition-level reward shaping via a learned reward estimator. Consistency regularization Bachman et al. (2014) to zero-reward transitions is adopted for the optimization of the reward estimator.
The rationale of SSRS is substantiated by identifying state-action pairs analogous to transitions with actual environmental rewards, and by conferring denser rewards accordingly upon those deemed sufficiently confident as “good” state-action pairs. Our empirical investigation reveals that the trajectory space conforms to the mild assumptions requisite for SSL, thereby permitting the application of SSL methodologies to refine the evaluative efficacy of the inferred reward function over zero-reward transitions.
The main contributions of the paper are listed as below:
  • We propose a novel reward shaping framework named SSRS, which operates on any base RL algorithm and employs SSL techniques to improve sample efficiency in sparse-reward environments.
  • We empirically examine the underlying conditions for SSL to be effective in RL settings, propose an additional data perturbation method to better process agent observations, and suggest potential directions for future applications of SSL in reinforcement learning.
  • We evaluate the performance of SSRS in reward-sparse Atari and robotic manipulation environments, comparing it with other supervised-based approaches—SORS Memarian et al. (2021), RCP Kumar et al. (2019)—and the standard RL baselines. Our framework demonstrates superior performance across the tested scenarios.
The remainder of this paper is structured as follows. Section. 2 introduces related work in reward shaping and SSL. Section. 3 presents the proposed framework alongside the necessary background on SSL. Section. 3 describes the experimental setup, provides comprehensive validation results, and offers an in-depth analysis of SSL.
2 Related Work
Sparse reward is a common challenge in RL. In addition to reward shaping, curriculum learning and hierarchical reinforcement learning are also effective approaches to addressing this issue. Curriculum learning structures training data through a progressive sequence, transitioning the agent from elementary to increasingly intricate tasks. Recent research Sayar et al. (2024) introduces a mechanism that generates adaptive learning objectives without domain expertise, promoting efficient learning in complex environments. Hierarchical reinforcement learning addresses the issue by decomposing a complex RL problem into subproblems, deploying specialized agents to manage distinct tiers of decision-making. Bukharin et al. (2025) propose a hierarchical reward modeling framework, which streamlines reward design by leveraging structural hierarchies or surrogate feedback to automatically evaluate trajectories and train reward models.
@@ 2 Related Work
Our work seeks to leverage reward shaping, which relies on densifying the reward function and therefore mitigate challenges in exploration or credit assignment induced by sparse rewards. There are different reward shaping methods for different objectives: Count-based methods Choi et al. (2019); Ostrovski et al. (2017); Bellemare et al. (2016) incentivize agents based on the rarity of states, while curiosity-driven exploration methods reward the agent’s exploratory behavior. ICM Pathak et al. (2017) also pursues rarity, but operationalizes it by defining curiosity as the error in the agent’s ability to predict the consequences of its own actions in a visual feature space. This curiosity signal serves as an intrinsic reward, driving the agent to explore novel states. Andrychowicz et al. (2017) propose Hindsight Experience Replay (HER) that enables learning from failures by treating unachieved goals as alternative goals for experience replay. Approaches using supervised learning for reward shaping don’t rely on specific objectives. The study Kumar et al. (2019) employs supervised learning techniques, viewing non-expert trajectories collected from sub-optimal policies as optimal supervision to do the reward shaping. Memarian et al. (2021) leverage self-supervised methods to extract signals from trajectories while simultaneously updating the policy.
@@
We extend supervised learning to a semi-supervised paradigm and introduce SSL techniques into the RL framework, specifically employing consistency regularization and pseudo-labeling Bachman et al. (2014) in the trajectory space. In the realm of SSL, pseudo-labeling employs the model itself to generate artificial labels for unlabeled data, while consistency regularization leverages unlabeled data by capitalizing on the invariance assumption—that the model should yield consistent predictions when presented with perturbed variants of identical inputs. Berthelot et al. (2019); Sohn et al. (2020) further improve this methodology on image classification tasks. We will formally introduce these SSL techniques in the later Background section and justify how our reward estimator is optimized through them.
3 Methodology
@@
Figure 1: The figure illustrates how the SSRS method leverages both non-zero reward transitions and sparse reward transitions for reward shaping through a semi-supervised learning approach incorporating strong-weak data augmentation. The augmentation techniques include cutout, smooth, and entropy augmentation, among others (see Table. 2 in the Appendix for details).
@@
In this section, we introduce the proposed method and its pipeline is graphically illustrated in Figure. 1, which consists of a reward estimator optimized through semi-supervised learning, and the data augmentation technique for transitions that serves as the perturbation required for semi-supervised learning. We will elaborate the key parts of the method in the following subsections and the detailed steps are shown in Algorithm 1.
Given a Markov Decision Process (MDP) with an original reward function $r(s,a)$, reward shaping refers to modifying this reward function—either by replacing it entirely or augmenting it with an auxiliary function—to produce a new reward function $R(s,a)$. In this paper, we focus on transition-level reward shaping. Specifically, for a one-step transition at timestep $t$ represented as $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, the shaped transition becomes $\langle s_{t},a_{t},\hat{r_{t}},s_{t+1}\rangle$, where $\hat{r_{t}}=R(s_{t},a_{t})$ denotes the newly assigned reward for that transition. Since our method is of a classification nature, that is to classify zero-reward transitions by matching them to these observed non-zero rewards and accordingly assigning a corresponding reward value. Let $Z=\{z_{i},\ i=1,\dots,N_{z}\}$ be the set of $N_{z}$ collected non-zero rewards, for a transition $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$ the reward distribution $p(\hat{r_{t}}=z|S=s_{t},A=a_{t})$ indicates the probability of assigning reward $z\in Z$ subject to the normalization condition $\sum_{z\in Z}{p}(\hat{r_{t}}=z|S=s_{t},A=a_{t})=1$.
3.1 Background
Consistency regularization and pseudo-labeling are key components in modern state-of-the-art semi-supervised learning methods, which have been primarily applied in classification tasks. To contextualize their use, we draw on the literature introduced earlier to describe these techniques. Let $p(r|s,a)$ denote the predicted reward distribution generated by the model given input $s$ and $a$. Consistency regularization relies on the assumption that the model’s output should remain stable under input perturbations, allowing the model to be optimized in a semi-supervised manner via the loss function
$\sum_{s,a}\left\|p(r|\tilde{s},\tilde{a})-p(r|\tilde{s},\tilde{a})\right\|_{2}^{2},$ (1)
where $\tilde{s}$ and $\tilde{a}$ are stochastically perturbed versions of the original state $s$ and action $a$, and $p$ denotes the model (its dependence on parameters is omitted). Because the perturbations differ, the two terms in the equation generally produce distinct values.
Pseudo-labeling leverages the model’s own high‑confidence predictions on unlabeled data. Hard labels are obtained via $\arg\max$ operation over the predicted distribution $\mathbf{q}=p(r|s,a)$—here we refer to $q$ as the confidence score. Only predictions whose maximum confidence exceeds a threshold $\lambda$ are retained and optimized using the loss:
$\mathds{1}(\max(\mathbf{q})\geq\lambda)\cdot\mathcal{H}(\mathbf{q},\arg\max(\mathbf{q})),$ (2)
where $\mathds{1}(\cdot)$ is the indicator function and $\mathcal{H}$ denotes the cross-entropy.
3.2 Semi-Supervised Reward Estimator
@@
The reward estimator is the very module that takes (augmented) transitions as input and outputs estimated rewards (see Figure. 1). During the training phase, the estimator is optimized via consistency regularization using the cross-entropy. In the prediction phase, the final estimated reward is generated through $\arg\max$ selection, choosing the $z$ value in the reward set $Z$ corresponding to the peak of the confidence score distribution. Reward shaping in this study is framed as a classification task, within which the SSRS framework is designed to build a reward function that integrates a state-action-value reward distribution function $R_{Q}(s,a)$ and a state-value reward distribution function $R_{V}(s)$, so as to select the optimal reward from the set $Z$. The state-value reward distribution function $R_{V}(s):\mathcal{S}\to\Delta(Z)$ defined as $R_{V}(s)={p}(\hat{r}=z|S=s)$, maps a state to a probability distribution over the reward set $Z$, where $\Delta(Z)$ is the probability simplex. Correspondingly, the state-action value reward distribution function $R_{Q}(s,a):\mathcal{S}\times\mathcal{A}\to\Delta(Z)$ is given by $R_{Q}(s,a)={p}(\hat{r}=z|S=s,A=a)$.
Given an one-step transition $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, the estimated reward can be calculated from the estimations of $R_{V}(s)$ and $R_{Q}(s,a)$ as follows. We first define the confidence score vector of timestep $t$ over the reward set $Z$ in Eq.(3). The elements of reward set $Z$ is the true reward value collected in the trajectories. The cardinality of this set $N_{z}$, is a predefined hyperparameter that controls the number of reward candidates. By the following Eq.(3), we can get confidence vector $\mathbf{q_{t}}\in\mathbb{R}^{N_{z}}$ at timestep $t$, with each element $q_{i}\in\mathbf{q_{t}}$ corresponding to a estimated reward $z_{i}\in Z$, computed as
$\mathbf{q_{t}}=\frac{1}{2}(R_{Q}(s_{t},a_{t})+R_{V}(s_{t+1})).$ (3)
The estimate reward value $z_{t}$ is selected with maximum confidence score above the threshold $\lambda$, i.e., $z_{t}=\alpha(\mathbf{q_{t}},\lambda)$, where
@@
Note that we will use $\mathbf{q}_{t}(\theta)$ in subsequent sections to explicitly denote its dependence on $\theta$, which consists of $\theta_{1}$ (the parameter of the $R_{Q}$ function) and $\theta_{2}$ (the parameter of the $R_{V}$ function).
@@
In order to obtain the optimal $\theta$ value for accurately estimating the reward value of each transition $\langle s_{t},a_{t},r_{t},s_{t+1}\rangle$, SSRS framework first optimizes the following loss function using only transitions with non-zero rewards, so as to minimize the discrepancy between the reward estimator’s predictions and the ground truth values for these transitions:
@@
$L_{r}=\frac{1}{B}\sum_{t=1}^{B}\mathds{1}(\max(\mathbf{q}_{t}(\theta))\geq\lambda)\left({r_{t}}-\alpha(\mathbf{q_{t}}(\theta),\lambda)\right)^{2},$ (4)
where $B$ is the batch size of transitions.
@@
SSRS integrates consistency regularization and pseudo-labeling, employing data augmentation in a manner similar to Sohn et al. (2020) to optimize the reward estimator on zero-reward transitions. In sparse-reward settings, applying augmentations to both zero- and non-zero reward transitions further enhances the generalization capability of the estimator. Explicitly, for transitions with zero rewards, the reward estimator computes the loss as shown in Eq.(5), which is the loss between the strong augmentation term and weak augmentation term following typical loss design in consistency regularization. Assuming continuity of trajectories in the metric space, it calculates the confidence of reward values after weak and strong augmentations, denoted by $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$, respectively. The calculation of $\mathbf{q_{t}^{w}}$ and $\mathbf{q_{t}^{s}}$ follows Eq.(3), simply replacing the input state s with its weakly or strongly augmented version. Here we omit their dependence on $\theta$ for simplicity. For confidence values greater than the threshold in weak augmentation, the one-hot operation is performed Bachman et al. (2014), denoted as $\mathbf{q_{t}^{w,\,\prime}}$, and then used to compute the cross-entropy loss with the normalized confidence values greater than the threshold in strong augmentation, denoted as $\mathbf{q_{t}^{s,\,\prime}}$. The loss function is denoted as.
$L_{s}=\frac{1}{B}\sum_{t=1}^{B}$ $[\mathds{1}(\max(\mathbf{q_{t}^{s}})\geq\lambda,\,\max(\mathbf{q_{t}^{w}})\geq\lambda)$
$\cdot\mathcal{H}(\mathbf{q_{t}^{w,\,\prime}},\mathbf{q_{t}^{s,\,\prime}}),$ (5)
@@
Weak and strong augmentations follow the specifications in Table. 2 in Appendix, where we consider Gaussian noise with smaller parameters as weak augmentation, and other augmentation such as smoothing, translation, and cutout, with larger parameters as strong augmentation. Unlike conventional image-based augmentation techniques (e.g., cutout, flipping, rotation), we exploit the unitless nature of entropy—which quantifies uncertainty in a probability distribution independently of scale or domain—to design a novel entropy augmentation method suitable for non-image tasks within SSRS. Implementation details are provided in the following section.
3.3 Extensions of SSRS
Since the estimator employs two neural networks to approximate $R_{V}(s)$ and $R_{Q}(s,a)$ respectively, we introduce a constraint to regulate their outputs and stabilize the training process. $R_{V}(s)$ provides a global baseline for a state, while advantage function captures the relative advantage or disadvantage of action $a$ with respect to this baseline. This separation makes it easier to assess the relative importance of specific actions, and thus compensates for the insensitivity to actions because the disturbance of data augmentation only takes place on state $s$. We introduce the monotonicity constraint in the SSRS framework as follows, which quantifies the mean square positive advantage function values. Define advantage function $\delta_{t}(\theta)=R_{Q}(s_{t},a_{t},\theta_{1})-R_{V}(s_{t},\theta_{2})$ where $\theta$ is the combination of $\theta_{1}$ and $\theta_{2}$, we have
$L_{QV}=\begin{cases}0,&\delta_{t}(\theta)>0,\\ \displaystyle\frac{1}{B}\sum_{t=1}^{B}\left(\delta_{t}(\theta)\right)^{2},&\text{else}.\end{cases}$ (6)
Minimizing the objective loss $L_{QV}$ over the parameter $\theta$ improves SSRS performance, which will be validated in experiment section.
Upon incorporating the aforementioned extensions, the algorithm for value-based SSRS with synchronous update —i.e., value function update and reward shaping carry out simultaneously— is outlined in Algorithm 1. Since the reward estimator does not distinguish between augmented and non-augmented transitions, the loss function computation can be parallelized, making the algorithm implementation quite simple and straightforward. The agent begins by collecting experience transitions. The reward estimation network is then initialized through back-propagation, which can be performed either synchronously or asynchronously using the loss function. The complete SSRS loss function, as specified in Eqs. (4), (6) and (5), combines these components through the consistency coefficient $\beta$. In Appendix, we provide the derivations of the “approximate” gradients of loss functions $L_{QV}$, $L_{r}$ and $L_{s}$ with respect to $\theta$ in Appendix, after approximating the indicator functions by sigmoid function.
3.4 Entropy Augmentation for SSRS
As a form of SSL perturbation, the implementation of entropy augmentation primarily utilizes Shannon Entropy for computation. Shannon Entropy Shannon (1948) quantifies the amount of information required to describe or encode a random variable’s possible states. In RL tasks, observations and trajectories are matrices in practice, but can also be normalized and viewed as discrete random variable. Consider a $m\times n$ matrix $A$ with each element $a_{ij}\geq 0,\,\forall i=1,...,m,\,j=1,...,n$, and the Shannon Entropy of matrix $A$ is defined by
$H(A):=-\sum_{k=1}^{mn}p_{k}\log(p_{k}),$ (7)
Denote state space as $\mathcal{S}\in\mathbb{R}^{m_{1}}$ and action space as $\mathcal{A}\in\mathbb{R}^{m_{2}}$, with $m_{1},\,m_{2}\geq 1$. Consider a trajectory $\tau$ of length $N$ as $\langle s_{0},a_{0},r_{0},s_{1},\allowbreak a_{1},r_{1},\dots,s_{N}\rangle$ with $a_{t}\in\mathcal{A},\ s_{t}\in\mathcal{S}$, for all $t=1,\dots,N$, and rewards are given following the environment’s original reward function. The set of all trajectories can be stacked into the matrix $\bar{\Gamma}=\begin{bmatrix}S,\,A,\,R\end{bmatrix}$ with $S\in\mathbb{R}^{N\times m_{1}},\,A\in\mathbb{R}^{N\times m_{2}}$ and $R\in\mathbb{R}^{N\times 1}$, and now we define the transform entropy augmentation $\sigma$ over matrix $\bar{\Gamma}$ as: